Liquid of specific gravity 0.90 flows in a horizontal tube 6.0 in diameter. In a section where the tube is constricted to 4.2 cm in diameter, the liquid pressure is less than that in the main tube by 16,000 dyn/cm
1
Expert's answer
2021-09-02T10:24:51-0400
"P=0.5\rho v^2\\1600=0.5(900)v^2\\v=1.9\frac{m}{s}"